# Statistical Inference with Limited Memory: A Survey

### v1

@@ 8 Distribution Property Testing
To obtain the $\Omega(k)$ lower bound, we focus first on the case of ergodic Markov chains. The proof is based on analysis of stationary distributions and properties of Markov chain transition matrices null space, and it shows that an input distribution can be perturbed in a way that leaves the stationary distribution of the chain unchanged. Thus, there exists a perturbation of the uniform distribution that results in a distribution $\varepsilon$ far from uniform, but induces the same stationary distribution on the state space of the chain. For a given FSM, denote the transition matrix and stationary vector induced by a distribution $p$ as $\operatorname{\mathbf{P}}(p)$ and $\pi_{p}$, respectively. We can write the stationary equations under the uniform distribution, $\pi_{U_{k}}=\pi_{U_{k}}\operatorname{\mathbf{P}}({U_{k}})$, as $\pi_{U_{k}}=MA{U_{k}}$, where $M\in[0,1]^{S\times S^{2}}$ is of the form
@@
and $A\in[0,1]^{S^{2}\times k}$ is a matrix whose rows are indexed by $(i,j)\in[S]\times[S]$, such that the entry $A_{(i,j),m}$ specifies the probability of moving to state $j$ from state $i$ when the input is $m\in[k]$. To simplify, we can write $\pi_{U_{k}}=MA{U_{k}}=B{U_{k}}$, where $B=MA\in[0,1]^{S\times k}$. If $S<k$, the matrix $B$ must have a nontrivial kernel of rank at least $k-S$. Letting $V$ be the subspace spanned by $\ker(B)\cap\ker([1\ldots 1])$, we have that $\dim(V)\geq k-1-S\triangleq K$. Using properties of norm contraction, one can show that there exists some vector $x\in V$ with $\lVert x\rVert_{\infty}=1/k$ and $\lVert x\rVert_{1}\geq K/k$. For this $x$, we can define $Q={U_{k}}-x$. Recalling that $V$ is orthogonal to $[1\ldots 1]$, we have that $Q$ is a valid probability distribution and $d_{\mathsf{TV}}{(U_{k},Q)}=\frac{1}{2}\lVert x\rVert_{1}\geq K/2k$. Since $x\in\ker(B)$, we thus have that $BQ=B{U_{k}}=\pi_{U_{k}}$. But $BQ=\pi_{U_{k}}$ are simply the stationary equations for the chain under $Q$, thus $\pi_{Q}=\pi_{U_{k}}$. This implies that for any finite state machine with $S=(1-2\varepsilon)k-1$ states, there exists a distribution $Q$ with $d_{\mathsf{TV}}{(U_{k},Q)}\geq\varepsilon$ such that $\pi_{U_{k}}=\pi_{Q}$, so the machine cannot distinguish between $U_{k}$ and $Q$. Thus, an ergodic finite state machine that attains non-trivial error must have $\Omega(k)$ states.
@@ 8.1 Additional Results
In a recent work, Canonne and Yang [126] purposed simpler algorithms for uniformity testing with $n$ samples that use $b$ bits of memory. Their aim was to simplify the involved analysis of [79] and relax some of the limitations on the range of parameters allowed. In their setup, the restriction on the memory size are $b\geq\max\{\log k,\log n\}$ and $b\leq\min\{n\log k,k\log n\}$. The first upper bound follows since straightforward lossless encoding takes $n\log k$ bits, and the second follows since only keeping the number of times each symbol is seen among the $n$ samples, which is a sufficient statistic for the problem, takes $k\log n$ bits. Their first (deterministic) algorithm achieves the same upper bound as [79] without the severe memory limitations of $b>>\frac{\log k}{\varepsilon^{2}}$. It relies on a uniformity testing algorithm purposed by Diakonikolas et al. [127], which is the total variation distance between the empirical distribution and the uniform distribution. Letting $N_{i}$ denote the frequency of the symbol $i$, and assuming that $n\leq k$, this statistic is just the normalized number of unseen symbols, that is
@@
The second (randomized) algorithm relies on the primitive of domain compression introduced by Acharya et al. [62], a variant of hashing tailored to distribution testing and learning, where one can transform an instance of testing over domain size $k$ and distance parameter $\varepsilon$ to a new instance over domain size $k^{\prime}$ and (smaller) distance parameter $\varepsilon^{\prime}\asymp\varepsilon\cdot\sqrt{\frac{k^{\prime}}{k}}$. Letting $k^{\prime}\asymp b/\log n$, the memory can now save all the ’new’ samples, and due to known results on uniformity testing, the resulting algorithm works as long as the number of samples $n$ satisfies
@@ 9.1 Entropy Estimation
$\operatorname{\mathsf{P_{e}}}(P,\hat{\pi})=\limsup_{n\rightarrow\infty}\Pr\left(|\hat{\pi}(T_{n})-H(P)|>\varepsilon\right),$ (46)
@@
We are generally interested in $\mathsf{MC}(\delta)$, i.e., the smallest number of states needed to guarantee $\varepsilon$-approximate error probability of at most $1/3$ for all $P\in\Delta_{k}$.
@@
In a very recent result [137], Berg et al. have proved that $\log k+2\log(1/\varepsilon)+o\left(\log k\right)$ bits suffice for entropy estimation when $\varepsilon>10^{-5}$, thus showing that the constant number of $\log k$ bits needed can be reduced all the way down to $1$ (barring that $\varepsilon$ is not too small). Furthermore, the proposed algorithm obtaining the upper bound was shown to converge within $\tilde{O}(k^{c})$ samples, for any $c>1$. The algorithm approximates the logarithm of $P(x)$, for a given $x\in[k]$, using a Morris counter  [138], a concept we now briefly explain: Suppose one wishes to implement a counter that counts up to $m$. Maintaining this counter exactly can be accomplished using $\log m$ bits. Morris gave a randomized “approximate counter”, which allows one to retrieve a constant multiplicative approximation to $m$ with high probability using only $O(\log\log m)$ bits. The Morris Counter was later analyzed in more detail by Flajolet [21], who showed that $O(\log\log m+\log(1/\varepsilon)+\log(1/\delta))$ bits of memory are sufficient to return a $(1\pm\varepsilon)$ approximation with success probability $1-\delta$ (this was later improved by Nelson and Yu [139], who showed that $O(\log\log m+\log(1/\varepsilon)+\log\log(1/\delta))$ bits suffice).
@@
The inherent structure of the Morris counter is particularly suited for constructing a nearly-unbiased estimator for $\log P(x)$, which makes it a judicious choice as a component in memory efficient entropy estimation. The algorithm of [137] uses two Morris counters, one that approximates a clock, and one that approximates $\log P(X)$. In order to compute the mean of these estimators, $\operatorname{\mathbb{E}}[\widehat{\log P(X)}]$, in a memory efficient manner, a finite-memory bias estimation machine (e.g., [110, 87]) is leveraged for simulating the expectation operator. The bias estimation machine is incremented whenever a count is concluded in the Morris counter approximating the clock, thus essentially averaging our estimates over all $x$ values. With $\tilde{O}(k^{c})$ samples, this averaging converges (approximately) to the mean of $-\log P(X)$, and thus outputs an approximation to the true underlying entropy.

### v2

@@ 8 Distribution Property Testing
To obtain the $\Omega(k)$ lower bound, consider first the case of ergodic Markov chains. The proof is based on analysis of stationary distributions and properties of Markov chain transition matrices null space, and it shows that an input distribution can be perturbed in a way that leaves the stationary distribution of the chain unchanged. Thus, there exists a perturbation of the uniform distribution that results in a distribution $\varepsilon$ far from uniform, but induces the same stationary distribution on the state space of the chain. For a given FSM, denote the transition matrix and stationary vector induced by a distribution $p$ as $\operatorname{\mathbf{P}}(p)$ and $\pi_{p}$, respectively. We can write the stationary equations under the uniform distribution, $\pi_{U_{k}}=\pi_{U_{k}}\operatorname{\mathbf{P}}({U_{k}})$, as $\pi_{U_{k}}=MA{U_{k}}$, where $M\in[0,1]^{S\times S^{2}}$ is of the form
@@
and $A\in[0,1]^{S^{2}\times k}$ is a matrix whose rows are indexed by $(i,j)\in[S]\times[S]$, such that the entry $A_{(i,j),m}$ specifies the probability of moving to state $j$ from state $i$ when the input is $m\in[k]$. To simplify, we can write $\pi_{U_{k}}=MA{U_{k}}=B{U_{k}}$, where $B=MA\in[0,1]^{S\times k}$. If $S<k$, the matrix $B$ must have a nontrivial kernel of rank at least $k-S$. Letting $V$ be the subspace spanned by $\ker(B)\cap\ker([1\ldots 1])$, we have that $\dim(V)\geq k-1-S\triangleq K$. Using properties of norm contraction, one can show that there exists some vector $x\in V$ with $\lVert x\rVert_{\infty}=1/k$ and $\lVert x\rVert_{1}\geq K/k$. For this $x$, we can define $Q={U_{k}}-x$. Recalling that $V$ is orthogonal to $[1\ldots 1]$, we have that $Q$ is a valid probability distribution and $d_{\mathsf{TV}}{(U_{k},Q)}=\frac{1}{2}\lVert x\rVert_{1}\geq K/2k$. Since $x\in\ker(B)$, we have that $BQ=B{U_{k}}=\pi_{U_{k}}$. But $BQ=\pi_{U_{k}}$ are simply the stationary equations for the chain under $Q$, thus $\pi_{Q}=\pi_{U_{k}}$. This implies that for any finite state machine with $S=(1-2\varepsilon)k-1$ states, there exists a distribution $Q$ with $d_{\mathsf{TV}}{(U_{k},Q)}\geq\varepsilon$ such that $\pi_{U_{k}}=\pi_{Q}$, so the machine cannot distinguish between $U_{k}$ and $Q$. Thus, an ergodic finite state machine that attains non-trivial error must have $\Omega(k)$ states.
@@ 8.1 Additional Results
In a recent work, Canonne and Yang [132] purposed simpler algorithms for uniformity testing with $n$ samples that use $b$ bits of memory. In their setup, the restriction on the memory size are $b\geq\max\{\log k,\log n\}$ and $b\leq\min\{n\log k,k\log n\}$. The first upper bound follows since straightforward lossless encoding takes $n\log k$ bits, and the second follows since only keeping the number of times each symbol is seen among the $n$ samples, which is a sufficient statistic for the problem, takes $k\log n$ bits. Their first (deterministic) algorithm achieves the same upper bound as [85], and relies on a uniformity testing algorithm purposed by Diakonikolas et al. [133], which is the total variation distance between the empirical distribution and the uniform distribution. Letting $N_{i}$ denote the frequency of the symbol $i$, and assuming that $n\leq k$, this statistic is just the normalized number of unseen symbols, that is,
@@
The second (randomized) algorithm relies on the primitive of domain compression introduced by Acharya et al. [71], a variant of hashing tailored to distribution testing and learning, where one can transform testing over domain size $k$ and distance parameter $\varepsilon$ to testing over domain size $k^{\prime}$ and (smaller) distance parameter $\varepsilon^{\prime}\asymp\varepsilon\cdot\sqrt{\frac{k^{\prime}}{k}}$. Letting $k^{\prime}\asymp b/\log n$, the memory can now save all the ’new’ samples, and due to known results on uniformity testing, the resulting algorithm works as long as the number of samples $n$ satisfies
@@ 9.1 Entropy Estimation
$\operatorname{\mathsf{P_{e}}}(P,\hat{\pi})=\limsup_{n\rightarrow\infty}\Pr\left(|\hat{\pi}(X^{n})-H(P)|>\varepsilon\right),$ (42)
@@
We are generally interested in $\mathsf{MC}(1/3)$, i.e., the smallest number of states needed to guarantee $\varepsilon$-approximate error probability of at most $1/3$ for all $P\in\Delta_{k}$.
@@
In a very recent result [143], Berg et al. proved that $\log k+2\log(1/\varepsilon)+o\left(\log k\right)$ bits suffice for entropy estimation when $\varepsilon>10^{-5}$, thus showing that the constant number of $\log k$ bits needed can be reduced all the way down to $1$ (barring that $\varepsilon$ is not too small). Furthermore, the proposed algorithm obtaining the upper bound was shown to converge within $\tilde{O}(k^{c})$ samples, for any $c>1$. The algorithm approximates the logarithm of $P(x)$, for a given $x\in[k]$, using a Morris counter  [144], a concept we now briefly explain: Suppose one wishes to implement a counter that counts up to $m$. Maintaining this counter exactly can be accomplished using $\log m$ bits. Morris gave a randomized “approximate counter”, which allows one to retrieve a constant multiplicative approximation to $m$ with high probability using only $O(\log\log m)$ bits. The Morris Counter was later analyzed in more detail by Flajolet [29], who showed that $O(\log\log m+\log(1/\varepsilon)+\log(1/\delta))$ bits of memory are sufficient to return a $(1\pm\varepsilon)$ approximation with success probability $1-\delta$ (this was later improved by Nelson and Yu [145], who showed that $O(\log\log m+\log(1/\varepsilon)+\log\log(1/\delta))$ bits suffice).
@@
The inherent structure of the Morris counter is particularly suited for constructing a nearly-unbiased estimator for $\log P(x)$, which makes it a judicious choice as a component in memory efficient entropy estimation. The algorithm of [143] uses two Morris counters, one that approximates a clock, and one that approximates $\log P(X)$. In order to compute the mean of these estimators, $\operatorname{\mathbb{E}}[\widehat{\log P(X)}]$, in a memory efficient manner, a finite-memory bias estimation machine (e.g., [115, 92]) is leveraged for simulating the expectation operator. The bias estimation machine is incremented whenever a count is concluded in the Morris counter approximating the clock, and it results in essentially an average of the estimates over all $x$ values. With $\tilde{O}(k^{c})$ samples, this averaging converges (approximately) to the mean of $-\log P(X)$, and thus outputs an approximation to the true underlying entropy.
10 Conclusions, Outlook, and Open Problems
We discussed the problem of statistical inference under memory constraints and surveyed the state-of-the-art in several problems, including binary hypothesis testing, bias estimation, uniformity testing and entropy estimation. We focused our attention on these particular problems since we see them as canonical problems in statistics, but essentially any other problem in statistical inference can be studied under limited memory constraints. We refrain from specifying particular examples of such open problems, with only one notable exception - the linear regression problem. While this is arguably one of the most extensively studied problems in statistics, current understanding of how this problem is affected by memory constraints remains quite limited. In Section 7.2 we described some known results in this context, and pointed out their limitations. Ideally, one would like to characterize the optimal dependence of the $\ell_{2}$ loss (or any other loss) on the number of samples $n$, the number of memory states $S$, the dimension $d$ and the noise variance $\sigma^{2}$, in the setup $Y_{i}=\theta^{T}X_{i}+V_{i}$, $i=1,\ldots,n$, where $\theta\in\mathbb{R}^{d}$ belongs to a bounded set or is sparse (or is drawn from some prior), $X_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{D}$ for some standard distribution $\mathcal{D}$, say $\mathcal{D}=\mathcal{N}(0,I)$, and $V_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,\sigma^{2})$. First order methods, such as stochastic gradient descent, often give a reasonable tradeoff between the memory and sample cost in this model, but the optimal loss remains unknown (see [120]).
We have seen that Markov chain theory plays a major role in characterizing the minimax risk in limited memory setting, often in deriving upper bounds by analyzing the induced limiting distributions for certain carefully designed finite-state machines. But this methodology was also used in [93] to obtain a lower bound in the uniformity testing problem, by showing that if a machine has a small number of states, then there must exist some distribution that is far from uniform but induces the same stationary distribution on the state space as the uniform distribution. This observation suggests that the problem of statistical inference under memory constraints, in the limit of many samples, is in fact equivalent to question of Markov chain instability, namely, to characterizing how sensitive a given chain topology can be to its driving distribution. Let us make this concrete. For a state space with $S$ states and a sample alphabet size $k$, a chain topology is a collection =$\mathcal{M}=\{M_{s}\in\{0,1\}^{k\times S}\}_{s\in[S]}$ of binary matrices, each satisfying $M_{s}\cdot\mathbf{1}=\mathbf{1}$. We interpret $(M_{s})_{i,s^{\prime}}=1$ as an edge connecting state $s$ to state $s^{\prime}$, when the input symbol is $i$. This formulation corresponds to deterministic $S$-state machines, while allowing $M_{s}\in\{[0,1]^{k\times S}\}_{s\in[S]}$ corresponds to randomized machines. Now, let $P$ be a probability distribution of the samples, represented as a column vector of length $k$. Define the function
$\xi_{S,\mathcal{M}}(P)=\frac{1}{S}\mathbf{1}^{T}\cdot\lim_{n\to\infty}\left((I_{S}\otimes P^{T})\cdot\mathsf{diag}\left(\{M_{s}\}_{s\in[S]}\right)\right)^{n},$ (45)
whenever the limit exists, where $\otimes$ is the Kronecker product, and $I_{S}$ is the $S$-dimensional identity matrix. In that case $\xi_{S,\mathcal{M}}(P)$ is the stationary distribution of an $S$-state chain with topology $\mathcal{M}$ under a driving distribution $P$. Now, we can phrase questions about the feasibility of limited-memory inference in terms of $\xi_{S,\mathcal{M}}(P)$. For example, in a simple binary hypothesis testing between $P$ and $Q$, we see that
$\sup_{\mathcal{M}}d_{\mathsf{TV}}\left(\xi_{\mathcal{S,M}}(P),\xi_{\mathcal{S,M}}(Q)\right)>1/2$ (46)
is a necessary condition for $S$ states to suffice. Or, for estimation in a parametric family $\{P_{\theta}\}$ under some metric loss function $\ell(\hat{\theta},\theta)$, we have that
$\sup_{\mathcal{M}}\inf_{\ell(\theta,\theta^{\prime})\geq\epsilon}d_{\mathsf{TV}}\left(\xi_{\mathcal{S,M}}(P_{\theta}),\xi_{\mathcal{S,M}}(P_{\theta^{\prime}})\right)>1/2$ (47)
is a necessary condition to attain a risk $\epsilon$. Hence, it is interesting to study the joint range
$\left\{\left(d_{\mathsf{TV}}\left(\xi_{\mathcal{S,M}}(P_{\theta}),\xi_{\mathcal{S,M}}(P_{\theta^{\prime}})\right),\ell(\theta,\theta^{\prime})\right)~{}:~{}\theta,\theta^{\prime}\in\Theta,~{}\mathcal{M}\right\}$ (48)
consisting of all pairs of stationary total variation distance and loss that can be attained by any (deterministic) $S$-state machine.
This survey has focused mainly on characterizing the memory complexity of the problems that were considered, while assuming we have access to an unlimited number of samples. Characterizing the optimal scaling of the risk as a function of both the number of samples $n$ and the number of memory states $S$ is a significantly harder problem, which remains open even for the simplest models one can think of, e.g., the bias estimation model. Information theoretic tools, mainly variants of the information bottleneck concept, were found useful for deriving lower bounds in these scenarios, as described, e.g., in Section 5. The idea is to show that with a small number of memory states, we cannot accurately predict the next sample $X_{t+1}$ from the memory state $U$ that depends only on $X^{t}$ (which implies that we do not learn much about $\theta$ from $U$). To that end, one analyzes the information bottleneck function
$\mathrm{IB}(\log S)=\max_{P_{U|X^{t}}:I(X^{t};U)\leq\log S}I(X_{t+1};U).$ (49)
In some cases, the restriction to the Markov chain $U-X^{t}-X_{t+1}$ is too weak to capture the sequential nature of the problem, where $U$ is actually obtained by $t$ consecutive updates. We believe that developing stronger techniques for upper bounding $I(U;X_{t+1})$ is a key step in improving the lower bounds on memory constrained inference.
Finally, we mention a problem in communication with memory constraints, in the same spirit as the problems surveyed in this article, which is worth studying. Consider the problem of communicating over $n$ channel uses of a discrete memoryless channel (DMC) $P_{Y|X}$, when the decoder has limited memory of $S$ states. The question is what is the largest achievable rate. Specifically, an $(n,R,S,\varepsilon)$-code for the DMC $P_{Y|X}$ is specified by an encoder $f:\left[\lceil 2^{nR}\rceil\right]\to\mathcal{X}^{n}$, a decoder update function $g:\mathcal{Y}\times[S]\to[S]$ and a decision function $d:\left[S\right]\to\left[\lceil 2^{nR}\rceil\right]$. For a message $W\sim\mathrm{Uniform}\left[\lceil 2^{nR}\rceil\right]$ the channel input is $X^{n}=f(W)$. The channel output $Y^{n}$ is generated by passing $X^{n}$ through $P^{\otimes n}_{Y|X}$. The update function is then applied recursively with $U_{0}=1$ and $U_{t}=g(Y_{t},U_{t-1})$, $t=1,\ldots,n$. Finally, the estimate for $W$ is $\hat{W}=d(U_{n})$, and the error probability is $\varepsilon=\Pr(W\neq\hat{W})$. We denote
$R^{*}(n,P_{Y|X},S,\varepsilon)$ $=\max\{R~{}:~{}\exists(n,R,S,\varepsilon)-\text{code}\}$ (50)
$C_{\varepsilon}(P_{Y|X},S)$ $=\sup_{n\in\mathbb{N}}R^{*}(n,P_{Y|X},S,\varepsilon)$ (51)
and the goal is to characterize this quantity. Note that for finite output alphabet channels $C_{\varepsilon}(P_{Y|X},S)$ is lower bounded by the finite-blocklength fundamental limit (without memory constraints) $R^{*}(\log{S}/\log|\mathcal{Y}|,P_{Y|X},\infty,\varepsilon)$, since we can use the $S$ memory states to save $n_{S}=\log{S}/\log{|\mathcal{Y}|}$ channel outputs. The interesting question here is how much better we can do, if at all, by using $n>n_{S}$.